Solve:
3^(x^2-3x)=81

Answer by RAY100(1637)

3^(x^2 -3x) =81=3^4
.
x^2 -3x =4,,,,,,,exponents must be equal
.
x^2 -3x -4 =0
.
(x-4)(x+1) =0
.
x=4,-1
.
check,,(4),,,3^((4)^2 -3(4)) =81
.
3^(16-12) =3^4 = 81,,,,ok
.
(-1),,,,,3^( (-1)^2 -3(-1))= 3^(1+3) = 3^4 =81,,,,ok